# Superconducting Electronic Device Response Linearity Calculation Method Using Discrete Fourier Analysis

### 27 Feb 2024 v1

@@ III Algorithm
The main idea of the method for determining the linearity from the calculated data is quite simple. The input data for the method is the output signal of the device $V_{\text{signal}}(F)$ (where $F$ is a parameter [^1] In real devices, this could be, i.g., an external magnetic flux.) calculated by some program (PSCAN [7, 8], WRSpice [9], etc.), and the result of applying this method should be the linearity of this signal in terms of SFDR.
@@
    Suppose we have a calculated signal $V_{\text{signal}}(F)$, which is a set of $N$ points of the form $\left(V_{\text{signal}}\right)_{n}(F_{n})$. In general, the $F_{N}$ may not be distributed evenly. However, for simplicity, we will assume that they were calculated with a constant step $\Delta F$. Note that this signal does not yet contain the time-dependence. This is the device’s response function and not the quasi-harmonic signal under investigation. The modification of the method in the event that the simulation results in time dependence $V_{\text{signal}}(t)$ will be considered in the A.
@@
    Using $\mathfrak{V}_{\text{signal}}(F)$, we can obtain the device’s output signal when a one-tone signal with a certain frequency $\omega$ is applied to its input: $\mathfrak{V}_{one-tone}(t)=\mathfrak{V}\left(A\sin\left(\omega t\right)+A_{0}\right)$, where $A$ and $A_{0}$ are some constants (amplitude and offset, respectively). The choice of these constants is determined entirely by the choice of the part of the response we wish to investigate. The choice of frequency (in relative units) depends on the number of points $2^{P}$ used for further Fourier analysis. As we have the flexibility to change each of these quantities, we can set the frequency arbitrarily (if we assume that the response does not depend on frequency; otherwise, we need to use the modified method described in the A). Note that $\mathfrak{V}_{one-tone}(t)$, defined in this way, is definitely continuous in $t$.
@@ IV.2 FFT accuracy
FFT methods are characterized by high accuracy. Moreover, when using accurate FFT methods (such as, i.g., the Cooley-Tukey FFT algorithm [11]), the introduced error depends only on the floating-point accuracy $\varepsilon_{\text{float}}$ and is approximately equal to $\varepsilon_{\text{float}}\sqrt{P}$. This error can be estimated to be $10^{-10}$. Thus, the introduced error is considered to be insignificant.
IV.3 The issue of the effect of input signal inaccuracy on the obtained linearity value
@@ Acknowledgements.
Appendix A
@@
Since the methods used remain the same (only their order has changed), all the reasoning about the accuracy of the method presented in section IV remains unchanged as well.

### 02 Mar 2024 v2

@@ III Algorithm
The main idea of the method for determining the linearity from the calculated data is quite simple. The input data for the method is the output signal of the device $V_{\text{signal}}(F)$ (where $F$ is a parameter [^1] In real devices, this could be, e.g., an external magnetic flux.) calculated by some program (PSCAN [7, 8], WRSpice [9], etc.), and the result of applying this method should be the linearity of this signal in terms of SFDR.
@@
    Suppose we have a calculated signal $V_{\text{signal}}(F)$, which is a set of $N$ points of the form $\left(V_{\text{signal}}\right)_{n}(F_{n})$. In general, the $F_{N}$ may not be distributed evenly. However, for simplicity, we will assume that they were calculated with a constant step $\Delta F$. Note that this signal does not yet contain the time-dependence. This is the device’s response function and not the quasi-harmonic signal under investigation. The modification of the method in the event that the simulation results in time dependence $V_{\text{signal}}(t)$ will be considered in the Appendix A.
@@
    Using $\mathfrak{V}_{\text{signal}}(F)$, we can obtain the device’s output signal when a one-tone signal with a certain frequency $\omega$ is applied to its input: $\mathfrak{V}_{one-tone}(t)=\mathfrak{V}\left(A\sin\left(\omega t\right)+A_{0}\right)$, where $A$ and $A_{0}$ are some constants (amplitude and offset, respectively). The choice of these constants is determined entirely by the choice of the part of the response we wish to investigate. The choice of frequency (in relative units) depends on the number of points $2^{P}$ used for further Fourier analysis. As we have the flexibility to change each of these quantities, we can set the frequency arbitrarily (if we assume that the response does not depend on frequency; otherwise, we need to use the modified method described in the Appendix A). Note that $\mathfrak{V}_{one-tone}(t)$, defined in this way, is definitely continuous in $t$.
@@ IV.2 FFT accuracy
FFT methods are characterized by high accuracy. Moreover, when using accurate FFT methods (such as, e.g., the Cooley-Tukey FFT algorithm [11]), the introduced error depends only on the floating-point accuracy $\varepsilon_{\text{float}}$ and is approximately equal to $\varepsilon_{\text{float}}\sqrt{P}$. This error can be estimated to be $10^{-10}$. Thus, the introduced error is considered to be insignificant.
IV.3 The issue of selecting FFT parameters
We need to choose three parameters when processing an FFT signal: the frequency $\omega$, the number of points $2^{P}$, and the time $T$. By converting from a discrete function to a continuous one, we can avoid the standard problems associated with FFT analysis and instead use a rectangular window. For proper consideration, we just need to formulate the following requirements for $T$, $2^{p}$, and $\omega$: an arbitrary integer number of cycles should fit within $T$, so $T$ should be equal to $M\times 2\pi/\omega$, where $M$ is an integer; and for harmonic detection purposes, it’s reasonable to require that $2^{P}\gg\left(5\text{ to }10\right)\times\omega$.
IV.4 The issue of the effect of input signal inaccuracy on the obtained linearity value
@@ Acknowledgements.
Appendix A Modification of the algorithm if a one-tone signal is present in the initial signal.
@@
Since the methods used are the same (only their order has changed), all the reasoning about the accuracy of the method presented in section IV remains unchanged as well, except for the considerations about selecting FFT parameters. A more complex consideration of using windows is beyond the scope of this appendix. However, taking into account the known frequency of the one-tone signal, we can always choose a time range $T$ such that an integer number of periods fit into it.